From a population of 700 plants, 500 plants are healthy, rest are unhealthy. For a sample size of 10 plants, calculate the probabilities that—(a) Exactly 3 are unhealthy (b) at least 3 plants are unhealthy (c) at most 3 plants are unhealthy (d) Exactly 3 are healthy (e) None healthy?

1
Expert's answer
2022-03-15T12:04:47-0400

Let X be the random variable representing the number of unhealthy plants.

The probability that a plant is unhealthy p=200/700 = 2/7

Since each plant can either be healthy or unhealthy, there are only two outcomes. Therefore, X follows a binomial distribution given by the formula:

P(X=x) =nCxpx(1-p)n-x

For x = 0,1,2...,n.

For n=10 and p=2/7=0.2857


a)

P(X=3) = 10C3(0.2857)3(0.7143)7

=0.2655

b)

P(X≥3) = P(3)+ P(4)+P(5)+P(6)+P(7)+P(8)+P(9)+P(10)

P(X=3) =10C3(0.2857)3(0.7143)7=0.2655

P(X=4) =10C4(0.2857)4(0.7143)6=0.1858

P(X=5) =10C5(0.2857)5(0.7143)5=0.0892

P(X=6) =10C6(0.2857)6(0.7143)4=0.0297

P(X=7)=10C7(0.2857)7(0.7143)3= 0.0068

P(X=8) =10C8(0.2857)8(0.7143)2= 0.0010

P(X=9) =10C9(0.2857)9(0.7143)1=0

P(X=0) =10C10(0.2857)10(0.7143)0=0


Therefore;

P(X≥3)=0.2655+0.1858+0.0892+0.0297+0.0068+0.0010

=0.5782


c)

P(X ≤ 3)= P(0)+P(1)+P(2)+P(3)

P(X=0)=10C0(0.2857)0(0.7143)10=0.0346

P(X=1)=10C1(0.2857)1(0.7143)9= 0.1383


P(X=2) =10C2(0.2857)2(0.7143)8= 0.2489


P(X=3) =10C3(0.2857)3(0.7143)7=0.2655


Therefore,

P(X≤3)=0.0346+0.1383+0.2489+0.2655

=0.6873



d)

Let the probability of a healthy plant p=5/7=0.7143

Let U be the random variable representing the number of healthy plants.

Then;

P(U=3) =10C3(0.7143)3(0.2857)7

=0.0068


e)

P(none is healthy) =P(all are healthy) P(X=10) =10C10(0.2857)10(0.7143)0

=0